If 17 textbooks cost 150.45 pounds, how much will 28 cost?
(Direct and Inverse Propotion Question- Student In year 10)

Answer by chibisan(131)

17books = 150.45pounds
1book = 150.45/17 pounds
28books = 150.45/17 x 28 =247.8pounds